# Calibrated Value-Aware Model Learning with Stochastic Environment Models

### v1

@@ 3.2 The (m,b)-VAML Family
$\quad=\mathbb{E}\Bigg{[}\biggl{|}\hat{V}\left(\hat{x}^{(m)}\right)-\left[[\mathcal{T}_{\mathcal{P}^{\pi}}^{b}V_{\mathrm{tar}}]\left(x^{(m)}\right)\right]_{\mathrm{sg}}\biggr{|}^{2}\Bigg{|}$
$\quad\quad\quad\hat{x}^{(m)}\sim\hat{p}^{m}(\cdot|x),x^{(m)}\sim\mathcal{P}^{m}(\cdot|x)\Bigg{]},$ (2)
where $V_{\mathrm{tar}}$ is a target network (Mnih et al., 2013). Note that samples from the real environment are used to approximate $[\mathcal{T}_{\mathcal{P}^{\pi}}^{k}\hat{V}]$. The loss function and the relation of the model and environments rollout are visualized in Figure 1.
@@ 4 Analysis of decision-aware losses in stochastic environments
The goal of learning in MBRL is to recover an (approximately) optimal model and to learn a correct value function. As we have shown, minimizing the IterVAML loss perfectly results in a model that leads to a correct Bellman Operator. However, in practice, the inner expectation of the IterVAML loss has to be replaced by a sampling-based approximation. In addition, $(m,b)$-VAML is used in the MuZero algorithm to update the value function directly in addition to the model. We show when this leads to learning correct models and value functions asymptotically. An overview of our conclusions can be found in Table 1. Proofs are found in Appendix A.
@@ 4.1 Calibration of surrogate loss functions
A prominent example of an minimum-uncalibrated loss is Bellman Residual Minimization (BRM), which results in the double sampling issue. While the target loss is minimized by the ground truth value function, BRM actually chooses a function that minimizes a variance term in addition to the Bellman error.
@@ 4.2 Model learning bias with stochastic models
When using samples from a stochastic model to compute the loss function in Equation 2, we are left with a variance error term that is closely related to the double-sampling problem
@@
We refer to the loss as Corrected VAML (CVAML).
@@
Let ^$\hat{\mathcal{L}}_{\mathrm{CVAML},m}^{k}=\hat{\mathcal{L}}^{k}_{\mathrm{IterVAML},m}-\widehat{\mathrm{Var}}_{\hat{p}}^{i,k}.$ The variance-corrected loss ^$\hat{\mathcal{L}}_{\mathrm{CVAML},m}^{k}(\hat{P},V|x,x^{(m)})$ is a calibrated surrogate loss for $\mathcal{L}_{\mathrm{IterVAML},m}^{\mathcal{P}^{\pi}}(\hat{P},V|x,x^{(m)})$.
Analogous to the IterVAML case, the sample-based $(m,b\geq 1)$-VAML loss is an uncalibrated loss for learning a model. To address this issue, we will introduce a corresponding variance correction term.
@@ 4.3 Value learning bias in stochastic models
We show that the MuZero loss therefore is not guaranteed to recover the correct value function, even when we have a perfect (stochastic) model. To formalize the issue, we compare the solution found with $(m,b)$-VAML with the regular TD loss function
$\mathcal{L}_{\mathrm{TD}}(V|V_{\mathrm{tar}},x^{(m)},x^{(m+1)},r^{(m)})$
$\quad=\left(V(x^{(m)})-\left[r^{(m)}+\gamma V_{\mathrm{tar}}(x^{(m+1)})\right]\right)^{2}.$ (4)
@@
    The set is complete, meaning that $[\mathcal{T}_{\mathcal{P}^{\pi}}V_{\mathrm{tar}}]\in\mathcal{V}$ for some target function $V$,
@@
The problem in the loss function again depends on the variance of the value function with regard to the model. To overcome this issue, we can introduce another variance correction term, similar to ^$\hat{\mathcal{L}}_{\mathrm{CVAML}}$. This retains the advantage that the MuZero loss is used for both model and value updates. However, as we average over the value prediction from the model, we only guarantee that
@@ 4.4 Discussion of theoretical results
While the problem resulting from the use of stochastic models is resolvable, the MuZero loss is still insufficient for learning per state value function with stochastic models. In addition, we observe issues with learning accurate value functions with the $(m\geq 1,b\geq 1)$-VAML loss even after correctly calibrating it. If the corrected loss is used to update the model, and a standard model-free or model-based value estimate is used as the value function learning target, the correct value function can be learned.
@@
To obtain a calibrated loss, we propose using ^$\hat{\mathcal{L}}_{\mathrm{CVAML}}$ to update the model and ^$\hat{\mathcal{L}}^{\hat{p}}_{0,j}$, or ^$\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{0,j}$, to update the value function.
@@ 5 Calibration impact on finite state MDPs
To test our findings, we run the $(m,b)$-VAML losses on small, finite state Garnet problems (Bhatnagar et al., 2007). Every problem is generated by sampling $m$ successor states for each state $x_{i}$, and parameterizing the transitions with weights $\omega_{i,j}$ sampled uniformly iid for all successor states $x_{j}$. The transition distribution is given by $p(x^{\prime}_{i}|x_{j})={e^{\omega_{i,j}/\tau}}/{\sum_{i}e^{\omega_{i,j}/\tau}}$. Varying the temperature $\tau$ we can interpolate between a deterministic transition and a $m$ uniform one.
The models are parameterized with two learnable matrices $\varphi\in\mathbb{R}^{n\times k}$ and $\psi\in\mathbb{R}^{n\times k}$ so that $\hat{p}_{k}(x_{i}|x_{j})=\mathrm{softmax}(\hat{\omega}_{i,j})=\mathrm{softmax}(\varphi_{i}^{\top}\psi_{j})$. By varying $k$ we create a low-rank constraint. As Farahmand et al. (2017) shows, $(m,b)$-VAML should be used when the model has insufficient capacity to represent the environment. As a baseline, we use the model $\hat{p}_{\mathrm{KL},k}\in\operatorname*{arg\,min}_{\hat{p}}\mathrm{KL}(p||\hat{p})$.
@@
As the stochasticity increases, we see an advantage for the calibrated losses. $(1,1)$-VAML is not able to achieve any good results even when the underlying environment is deterministic. This corroborates the theoretical finding that the error depends on the model’s stochasticity and not on the environment. As we initialize the models with high entropy, $(1,1)$-VAML is unable to learn a high variance value.
Calibrated $(1,1)$-VAML still struggles in environments with low stochasticity. We find that due to the variance term over value functions, the model is prone to get stuck in local minima as it reduces the value function difference per state faster than the model predictive variance. This suggests that a $(m\geq 1,0)$-CVAML loss is preferable with stochastic environments.
@@ 6 Stochasticity and auxiliary models
In general, deterministic function approximations cannot capture the full transition distribution in stochastic environments. However, it is an open question whether a deterministic latent model can be sufficient for learning a value-equivalent model, as conjectured by Oh et al. (2017).
@@ 6.1 Deterministic models for stochastic environments
Showing the existence of a deterministic value-aware model relies on the continuity of the transition kernel and involved functions $\varphi$ and $V$. These are standard assumptions that are necessary to prove the existence and measurablity of standard functions such as the value (Bertsekas & Shreve, 1978).
@@
$f^{*}\in\operatorname*{arg\,min}_{\hat{f}}\mathbb{E}_{\rho,\mathcal{P}^{\pi}}\left[\hat{\mathcal{L}}_{\mathrm{IterVAML},1}(\hat{f},V_{\mathcal{X}}|V_{\mathcal{X}},x)\right].$
We can conclude that given a sufficiently flexible function class $\mathcal{F}$, $(m,b)$-VAML can recover an optimal deterministic model for value function prediction. Note that our conditions solely ensure the existence of a measurable function; the learning problem might still be very challenging.
Note that the conditions for our proof here are slightly different than those for our definition of an uncalibrated loss. Here we show that given a flexible enough function class, a deterministic model can sufficiently minimize the IterVAML loss. In the previous section, we showed that if the function class admits a perfect
@@ 6.2 Auxiliary losses for latent-space models
Most prior works use some form of Bootstrap-your-own-latent (BYOL) (Grill et al., 2020) loss, which minimizes the difference between the next states encoding $\varphi(x^{(m)})$ and the model prediction $f^{m}(x^{(0)})$. In the case of linear features, models, and value functions, recent works have shown that such a loss can greatly aid in learning meaningful features for value function prediction (Lyle et al., 2021; Tang et al., 2023; Ni et al., 2024; Voelcker et al., 2024). When the difference is computed with an $L_{2}$, the model loss is
^$\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{\mathrm{model},m}\left(\hat{f},\varphi,\hat{V}|x\right)$
=^$~{}=\underbrace{\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{m,b}\left(\hat{f},\varphi,\hat{V}|x\right)}_{(m,b)\mathrm{-VAML}}+\underbrace{\biggl{|}\hat{f}^{m}(\varphi(x))-\bigl{[}\varphi(x^{(m)})\bigr{]}_{\mathrm{sg}}\biggr{|}^{2}}_{\mathrm{auxiliary}}.$ (6)
@@
This is the case if and only if the value function is linear in the embedding features $\varphi$, which is also referred to as a linear expectation model (Wan et al., 2019). However, learning an embedding in which the value function is linear can be difficult and may not lead to stable model predictions in complex, high-dimensional environments.
@@ 7 Experiments with latent-space models
where $a^{(0:m-1)}$ is a sequence of actions of length $m$ starting from state $x^{(0)}$.
In the case of deterministic models, the auxiliary loss is simply the MSE between the next states’ representations and the model’s predictions:
$\quad=\left\|\hat{p}^{m}\left(\cdot\Big{|}\varphi(x),a^{(0:m-1)}\right)-\varphi(x^{(m)}\right\|^{2}.$ (10)
For training the policy and value function, we follow the data-mixing protocol suggested by Voelcker et al. (2025). We use both model-generated on-policy and real environment data from the replay buffer to train a value and policy head using the $(m,b)$-VAML loss with a twinned critic parametrization similar to TD3 (Fujimoto et al., 2018b). At every timestep, half the data in the minibatch is model simulated, the other half ground-truth environment data. This means the model influences the value function both through the encoder and through on-policy model-generated data.
@@
Results: We plot aggregated final performance over 20 random seeds per environment-loss pair with 95% CI, estimated with stratified percentile bootstrap (Patterson et al., 2024). Aggregations of final performance over several environments are visualized using the RLiable library (Agarwal et al., 2021).
@@
Finally, we observe an advantage of $(1,0)$-updates over the $(1,1)$-updates. While previous work has found IterVAML to be unstable (Lovatto et al., 2020; Voelcker et al., 2022), comparisons were not made with SOTA model architectures, latent models, and auxiliary tasks. As additional experiments (see Figure 8) show, the IterVAML loss is stable even without additional auxiliary losses, although adding the BYOL loss is necessary to achieve non-trivial performance in humanoid tasks.
@@
Figure 4: Sample efficiency curve for both dog (left) and humanoid (right). Per-task normalized return is aggregated over 30 seeds per environment and 3 tasks for the humanoid domain, and 4 – for the dog domain, with 95% bootstrapped confidence intervals shaded. In addition to a higher final return, we observe significantly earlier learning for the $(0,1)$-(C)VAML on the dog task.
@@ 8 Related work
Other decision-aware algorithms:  Several other works propose decision-aware model learning algorithms that do not directly minimize a value function difference. D’Oro et al. (2020) weigh the samples used for model learning by their impact on the policy gradient. Nikishin et al. (2021) uses implicit differentiation to obtain a loss for the model function with regard to the policy performance measure. To achieve the same goal, Eysenbach et al. (2022) and Ghugare et al. (2023) choose a variational formulation. Modhe et al. (2021) proposes to compute the advantage function resulting from different models instead of using the value function. Ayoub et al. (2020) presents an algorithm for on selecting models based on their ability to predict value function estimates and provide regret bounds with this algorithm.
Learning with suboptimal models:  Several works have focused on the broader goal of using models with errors without addressing the loss functions of the model. Among these, some attempt to correct models using information obtained during exploration (Joseph et al., 2013; Talvitie, 2017; Modi et al., 2020; Rakhsha et al., 2022, 2024), or to limit interaction with wrong models (Buckman et al., 2018; Janner et al., 2019; Abbas et al., 2020). Several of these techniques can be applied together with corrected $(m,b)$-VAML to improve the model and value function learning further. Finally, we do not focus on exploration, but Guo et al. (2022) show how auxiliary losses can be used not only to stabilize learning but also to improve exploration.

### v2

# Calibrated Value-Aware Model Learning with Probabilistic Environment Models

@@ 3.2 The (m,b)-VAML Family
$\quad=\mathbb{E}\Bigg{[}\biggl{|}\hat{V}\left(\hat{x}^{(m)}\right)-\left[[\mathcal{T}_{\mathcal{P}^{\pi}}^{b}V_{\mathrm{tar}}]\left(x^{(m)}\right)\right]_{\mathrm{sg}}\biggr{|}^{2}\Bigg{|}x\Bigg{]}$ (2)
where $V_{\mathrm{tar}}$ is a target network (Mnih et al., 2013) and $x^{(m)}$ and $\hat{x}^{(m)}$ are sampled independently from $\mathcal{P}^{\pi}$ and $\hat{p}$ respectively. Note that samples from the real environment are used to approximate $[\mathcal{T}_{\mathcal{P}^{\pi}}^{b}V_{\mathrm{tar}}]$. The loss function and the relation of the model and environments rollout are visualized in Figure 1.
@@ 4 Analysis of decision-aware losses in stochastic environments
The goal of learning in MBRL is to recover an (approximately) optimal model and to learn a correct value function. As we have shown, minimizing the IterVAML loss perfectly results in a model that leads to a correct Bellman Operator. However, in practice, the inner expectation of the IterVAML loss has to be replaced by a sampling-based approximation. In addition, $(m,b)$-VAML is used in the MuZero algorithm to update the value function directly in addition to the model. We show when this leads to learning correct models and value functions asymptotically. An overview of our conclusions can be found in Table 1.
@@ 4.1 Calibration of surrogate loss functions
An example of minimum-uncalibrated-ness is the double sampling issue in Bellman Residual Minimization (BRM). While the target loss is minimized by the ground truth value function, BRM actually chooses a function that minimizes the Bellman error and and additional variance term.
@@ 4.2 Model learning bias with stochastic models
Proofs for this section can be found in subsubsection A.2.2. When using samples from a stochastic model to compute the loss function in Equation 2, we are left with a variance error term that is closely related to the double-sampling problem
@@
We refer to the loss as Calibrated VAML (CVAML).
@@
The variance-corrected loss ^$\hat{\mathcal{L}}_{\mathrm{CVAML},m}^{k}(\hat{P},V|x,x^{(m)})$ is a calibrated surrogate loss for $\mathcal{L}_{\mathrm{IterVAML},m}^{\mathcal{P}^{\pi}}(\hat{P},V|x,x^{(m)})$.
Analogous to the IterVAML case, the sample-based $(m,b\geq 1)$-VAML loss is an uncalibrated loss for learning a model.
@@ 4.3 Value learning bias in stochastic models
We show that the MuZero loss therefore is not guaranteed to recover the correct value function, even when we have a perfect (stochastic) model. To formalize the issue, we compare the solution found with $(m,b)$-VAML with the regular TD loss function. Note that for this loss, we assume that $x_{(m)}$ is a fixed sample from the environment, not a random variable like $x^{(m)}$, with $x_{(m+1)}$ being its successor state in the ground truth environment. This distinction is important, as it is exactly what leads to the bias in the $(m,b\geq 1)$ loss.
$\mathcal{L}_{\mathrm{TD}}(\hat{V}|V_{\mathrm{tar}},x^{(m)},x^{(m+1)},r^{(m)})$
$\quad=\left(\hat{V}(x^{(m)})-\left[r^{(m)}+\gamma V_{\mathrm{tar}}(x^{(m+1)})\right]\right)^{2}.$ (4)
To drive home that this problem is independent of the model error, we show that the $(m,b\geq 1)$-VAML loss is uncalibrated even if we substitute the ground truth environment for the learned model, and focus solely on the value learning component.
@@
    The set is complete, meaning that $[\mathcal{T}_{\mathcal{P}^{\pi}}V_{\mathrm{tar}}]\in\mathcal{V}$ for some target function $V_{\mathrm{tar}}$,
@@
The problem in the loss function again depends on the variance of the value function with regard to the model. Proofs for this section can be found in subsubsection A.2.3.
To overcome this issue, we can introduce another variance correction term, similar to ^$\hat{\mathcal{L}}_{\mathrm{CVAML}}$. This retains the advantage that the MuZero loss is used for both model and value updates. However, as we average over the value prediction from the model, we only guarantee that
@@ 4.4 Discussion of theoretical results
While the problem resulting from the use of stochastic models is resolvable, the MuZero loss is still insufficient for learning per state value function with stochastic models. We observe issues with learning accurate value functions with the $(m\geq 1,b\geq 1)$-VAML loss even after correctly calibrating it. If the calibrated loss is used to update the model, and a standard model-free or model-based value estimate is used as the value function learning target, the correct value function can be learned.
@@
To obtain a calibrated loss, we propose using ^$\hat{\mathcal{L}}_{\mathrm{CVAML}}$ to update the model and a regular model-based or model-free TD loss to update the value function.
@@ 5 Calibration impact on finite state MDPs
To test our findings, we run the $(m,b)$-VAML losses on small, finite state Garnet problems (Bhatnagar et al., 2007). We use the MuZero-style $(1,1)$-VAML loss to update both model and value function, while the IterVAML-style $(1,0)$-VAML loss is only used to train the model, and we use a regular model-based TD loss to update the value function. For the baseline, we similarly use a model-based TD loss together with a KL-based loss.
We use $n$ to denote the size of the state space, $k$ for the number of successor states in the garnet, and $j$ for the rank of the model. Every problem is generated by sampling $k$ successor states for each state $x_{i}$, and parameterizing the transitions with weights $\omega_{i,j}$ sampled iid from a standard normal distribution for all successor states $x_{j}$. The transition distribution is given by $p(x^{\prime}_{i}|x_{l})={e^{\omega_{i,l}/\tau}}/{\sum_{i}e^{\omega_{i,l}/\tau}}$. Varying the temperature $\tau$ we can interpolate between a deterministic transition and a uniform one.
The models are parameterized with two learnable matrices $\varphi\in\mathbb{R}^{j\times n}$ and $\psi\in\mathbb{R}^{j\times n}$ so that $\hat{p}_{k}(x_{i}|x_{l})=\mathrm{softmax}(\hat{\omega}_{i,l})=\mathrm{softmax}(\varphi_{i}^{\top}\psi_{l})$. By varying $j$ we create a low-rank constraint. As Farahmand et al. (2017) shows, $(m,b)$-VAML should be used when the model has insufficient capacity to represent the environment. As a baseline, we use the model $\hat{p}_{\mathrm{KL},k}\in\operatorname*{arg\,min}_{\hat{p}}\mathrm{KL}(p||\hat{p})$, which we approximate with gradient descent. The matrices $\varphi$ and $\psi$ are initialized by drawing weights randomly from a normal distribution with a very small standard deviation.
@@
As the stochasticity increases, we see an advantage for calibrated losses. $(1,1)$-VAML is not able to achieve good results even in a deterministic environment. As we initialize the models with high entropy, $(1,1)$-VAML is unable to learn a correct value function. This corroborates the theoretical finding that the error depends on the model’s stochasticity and not on the environment.
Calibrated $(1,1)$-VAML still struggles in environments with low stochasticity. We find that the model is prone to get stuck in local plateaus as it reduces the value function difference per state faster than the model predictive variance. This suggests that a $(m\geq 1,0)$-CVAML loss is preferable with stochastic environments.
@@ 6 Stochasticity and auxiliary models
In general, deterministic function approximations cannot capture the full transition distribution in stochastic environments. However, it is an open question whether a deterministic latent model can be sufficient for learning a value-equivalent model, as conjectured by Oh et al. (2017). We settle this question now.
@@ 6.1 Deterministic models for stochastic environments
Showing the existence of a deterministic value-aware model relies on the continuity of the transition kernel and involved functions $\varphi$ and $V$. These are standard assumptions that are necessary to prove the existence and measurability of standard functions such as the value function (Bertsekas & Shreve, 1978).
@@
$f^{*}\in\operatorname*{arg\,min}_{\hat{f}}\mathbb{E}_{\rho,\mathcal{P}^{\pi}}\left[\mathcal{L}_{\mathrm{IterVAML},1}(\hat{f},V_{\mathcal{X}}|V_{\mathcal{X}},x)\right].$
Proofs for this section can be found in Subsection A.3.
We can conclude that given a sufficiently flexible function class $\mathcal{F}$, $(1,b)$-VAML can recover an optimal deterministic model for value function prediction. Note that our conditions solely ensure the existence of a measurable function; the learning problem might still be very challenging.
Note that the conditions for our proof here are slightly different than those for our definition of an uncalibrated loss. Here we show that given a flexible enough function class, a deterministic model can sufficiently minimize the IterVAML loss. In the previous section, we showed that if the function class admits a perfect stochastic model, but not a perfect deterministic one, the $(m,b)$-VAML loss will incur a calibration error. Here we show the conditions under which a perfect deterministic model exists.
@@ 6.2 Auxiliary losses for latent-space models
Most prior works use some form of Bootstrap-your-own-latent (BYOL) (Grill et al., 2020) loss, which minimizes the difference between the next states encoding $\varphi(x^{(m)})$ and the model prediction $f^{m}(x^{(0)})$. In the following we write $(\hat{f},\varphi)=\hat{p}$ when we want to highlight the different components of the learned model.
In the case of linear features, models, and value functions, recent works have shown that such a loss can greatly aid in learning meaningful features for value function prediction (Lyle et al., 2021; Tang et al., 2023; Ni et al., 2024; Voelcker et al., 2024). When the difference is computed with an $L_{2}$, the auxiliary loss is
^$\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{\mathrm{aux},m}\left((\hat{f},\varphi)|x\right)=\biggl{\|}\hat{f}^{m}(\varphi(x))-\bigl{[}\varphi(x^{(m)})\bigr{]}_{\mathrm{sg}}\biggr{\|}^{2}.$ (6)
@@
This is the case if and only if the value function is affine in the embedding features $\varphi$, which is also referred to as a linear expectation model (Wan et al., 2019). However, learning an embedding in which the value function is linear can be difficult and may not lead to stable model predictions in complex, high-dimensional environments.
@@ 7 Experiments with latent-space models
where $a^{(0:m-1)}$ is a sequence of actions of length $m$ starting from state $x^{(0)}$. In the case of deterministic models, the auxiliary loss is simply the MSE version introduced in Equation 7.
The full model loss used is
^$\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{\mathrm{model},m}\left((\hat{f},\varphi),\hat{V}|x\right)$
=^$~{}=\underbrace{\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{m,b}\left((\hat{f},\varphi),\hat{V}|x\right)}_{(m,b)\mathrm{-VAML}}+\underbrace{\hat{\mathcal{L}}^{\mathcal{P}^{\pi}}_{\mathrm{aux},m}\left((\hat{f},\varphi)|x\right)}_{\mathrm{auxiliary}},$ (10)
replacing $\hat{f}$ with $\hat{p}$ for the stochastic model version.
For training the policy and value function, we follow the data-mixing protocol suggested by Voelcker et al. (2025). We use both model-generated on-policy and real environment data from the replay buffer to train a value and policy head using the $(m,b)$-VAML loss with a twinned critic parametrization similar to TD3 (Fujimoto et al., 2018b). At every timestep, half of the next states in the minibatch are replaced with model-simulated states resulting from on-policy actions, the other half ground-truth environment data. This means the model influences the value function both through the encoder and through on-policy model-generated data.
@@
Results: We plot aggregated performance over 20 random seeds with 95% CI, estimated with stratified percentile bootstrap (Patterson et al., 2024). Aggregations of final performance over several environments are visualized using the RLiable library (Agarwal et al., 2021).
@@
Finally, we observe an advantage of $(1,0)$-updates over the $(1,1)$-updates. While previous work claimed that IterVAML is unstable (Lovatto et al., 2020; Voelcker et al., 2022), comparisons were not made with SOTA model architectures and auxiliary tasks. As additional experiments (see Figure 8) show, the IterVAML loss is stable even without additional auxiliary losses, although adding the BYOL loss is necessary to achieve non-trivial performance in humanoid tasks.
@@
Figure 4: Sample efficiency curve for both dog (left) and humanoid (right). Per-task normalized return is aggregated over 30 seeds per environment and 3 tasks for the humanoid domain, and 4 – for the dog domain, with 95% bootstrapped confidence intervals shaded. In addition to a higher final return, we observe significantly earlier learning for the $(1,0)$-(C)VAML on the dog task.
@@ 8 Related work
Other decision-aware algorithms:  Several other works propose decision-aware model learning algorithms that do not directly minimize a value function difference. D’Oro et al. (2020) weigh the samples used for model learning by their impact on the policy gradient. Nikishin et al. (2021) uses implicit differentiation to obtain a loss for the model function with regard to the policy performance measure. To achieve the same goal, Eysenbach et al. (2022) and Ghugare et al. (2023) choose a variational formulation. Modhe et al. (2021) proposes to compute the advantage function resulting from different models instead of using the value function. Ayoub et al. (2020) presents an algorithm for selecting models based on their ability to predict value function estimates and provide regret bounds with this algorithm.
Learning with suboptimal models:  Several works have focused on the broader goal of using models with errors without addressing the loss functions of the model. Among these, some attempt to correct models using information obtained during exploration (Joseph et al., 2013; Talvitie, 2017; Modi et al., 2020; Rakhsha et al., 2022, 2024), or to limit interaction with wrong models (Buckman et al., 2018; Janner et al., 2019; Abbas et al., 2020). Several of these techniques can be applied together with $(m,b)$-CVAML to improve the model and value function learning further. Finally, we do not focus on exploration, but Guo et al. (2022) show how auxiliary losses can be used not only to stabilize learning but also to improve exploration.